# On Predictability of Reinforcement Learning Dynamics for Large Language Models

### v2

@@ Abstract
Recent advances in reasoning capabilities of large language models (LLMs) are largely driven by reinforcement learning (RL), yet the underlying parameter dynamics during RL training remain poorly understood. This work identifies two fundamental properties of RL-induced parameter updates in LLMs: (1) Rank-1 Dominance, where the top singular subspace of the parameter update matrix nearly fully determines reasoning improvements, recovering over 99% of performance gains; and (2) Rank-1 Linear Dynamics, where this dominant subspace evolves linearly throughout training, enabling accurate prediction from early checkpoints. Extensive experiments across 8 LLMs and 7 algorithms validate the generalizability of these properties. More importantly, based on these findings, we propose AlphaRL, a plug-in acceleration framework that extrapolates the final parameter update using a short early training window, achieving up to 2.5× speedup while retaining >96% of reasoning performance without extra modules or hyperparameter tuning. This positions our finding as a versatile and practical tool for large-scale RL, opening a path toward principled, interpretable, and efficient training paradigm for LLMs. Our model and code will be available at: https://github.com/caiyuchen-ustc/Alpha-RL.
@@ 1 Introduction
While effective, these researches mainly focus on post-hoc interpretability (i.e., explaining the endpoints of training), leaving the RL process itself less explored (Wang et al., 2025b; Zhang et al., 2025; Feng et al., 2025). Understanding parameter dynamics during RL is essential: not only for optimizing RL paradigms, but also for shedding light on the emergence of reasoning capabilities. Hence, this work aims to reveal the black-box of RL process by addressing two fundamental questions: Do RL-guided parameter updates follow consistent patterns, and how do these patterns give rise to reasoning capabilities?
@@
To solve these, we conduct a step-wise analysis of the parameter update matrix $\Delta\bm{W}$ (i.e., the parameter difference between the RL-trained model and the base model). After applying mathematical tools such as orthogonal subspace projection (Cai & Cao, 2024), we uncover a striking phenomenon: performing singular value decomposition (SVD) (Koren et al., 2009) on $\Delta\bm{W}$ reveals that the top singular subspace, defined as the Rank-1 Subspace, almost entirely determines the reasoning gains from RL. That is, adding only the Rank-1 component of $\Delta\bm{W}$ to the base model is sufficient to recover nearly all of the RL-trained model’s reasoning improvements. More remarkably, this property holds not just at convergence, but at any intermediate step of RL training, as shown in Figure 2 (a). We formalize this finding as Property 1 (Rank-1 Dominance): Rank-1 Subspace of $\Delta\bm{W}$ determines the reasoning gains of LLMs throughout the RL training process.
@@
This property inspired us to probe how the Rank-1 Subspace evolves during RL training. By applying partial least squares (PLS) to track the dimension-wise trajectory of the Rank-1 Subspace across training steps, we observe an almost strictly linear upward trend, with the linear rate metric $R^{2}$ (Geladi & Kowalski, 1986) exceeding 0.96. Consequently, the Rank-1 Subspace at a target step can be accurately forecast from a short early window of training. We formalize this as Property 2 (Rank-1 Linear Dynamics): Rank-1 Subspace evolves approximately linearly with RL training process, yielding high predictability from early-stage checkpoints of RL process.
@@
To validate the generality, we conducted extensive experiments across a diverse set of 8 models (ranging from 7B to 32B parameters, e.g., Qwen2.5 (Qwen et al., 2025) and Qwen3 (Yang et al., 2025)) trained with 7 distinct advanced training algorithms (e.g., RLOO (Ahmadian et al., 2024), GRPO (DeepSeek-AI et al., 2025) and DAPO (Yu et al., 2025)). Our analysis shows that, for Property 1, the Rank-1 Subspace alone recovers an average of 99.17% of the reasoning capability. For Property 2, (1) the linearity of Rank-1 Subspace’s evolution exhibits an average $R^{2}$ of 0.914, and (2) predictions of its later state based on early-stage states achieve an average error of less than 5%. Crucially, control experiments with alternative training paradigms like supervised fine-tuning and distillation on the same models yielded neither property, demonstrating that these phenomena are distinctive characteristics of the RL process for LLMs. Detailed experimental setups and results are presented in Sections 2 and 3.
@@ 2 Dominance of Rank-1 Subspace (Property 1)
In this section, our objective is to analyze the effect of the Rank-1 Subspace of the parameter update matrix $\Delta W$ on reasoning gains. In Section 2.1, we first introduce the method for quantifying the contribution of the Rank-1 Subspace. Then, we exhibit experimental setups and main results in Section 2.2. In Section 2.3, we investigate the underlying causes of of Rank-1 dominance by exploring the norm of $\Delta W$, the rank of $\Delta W$, and model embeddings.
@@ 2.1 Rank-1 Subspace
We first describe the method for quantifying the contribution of the Rank-1 Subspace of parameter updates $\Delta\bm{W}$ to reasoning gains. Specifically, performing SVD on $\Delta\bm{W}$, we have:
@@ 2.2 Rank-1 Subspace & Reasoning gains
Experiment Seting. Our experiments are conducted on eight models, covering five RL algorithms, including PPO (Schulman et al., 2017), RLOO (Ahmadian et al., 2024), GRPO (DeepSeek-AI et al., 2025), Dr.GRPO (Liu et al., 2025), and DAPO (Yu et al., 2025), as well as Distillation (DIST) (Hinton et al., 2015) and Supervised Fine-tuning (SFT) (Ye et al., 2025). We comprehensively evaluate the reasoning performance of these models on six reasoning benchmarks, including AIME24, AIME25 (Ye et al., 2025) and MATH-500 (Lightman et al., 2023), to verify the robustness and generality of our findings. More detailed settings are provided in Appendix A.
@@ 2.3 Analysis of Rank-1 Dominance
Approximate Low-rank of $\Delta\bm{W}$. It is worth mentioning that, in the above experiments, we discovered a universal approximate low-rank (Zhang, 2015) property of $\Delta\bm{W}$ in RL, which is completely absent in SFT and DIST. Due to space limitations, we provide a detailed discussion of this phenomenon in Appendix B. Furthermore, we propose that the superior properties observed in RL-trained LLMs (e.g., minimizing catastrophic forgetting and improving generalization) may fundamentally arise from this low-rank structure, which plays a pivotal role in the model’s ability to effectively retain and adapt learned knowledge. Additionally, we also observed in our experiments the unique role of Rank-1 in guiding the reasoning chain, where modifying a small number of tokens achieves reasoning performance comparable to that of the fully trained model. We recommend interested readers to refer to the detailed results and analysis in Appendix C.
@@ 3.1 Exploring the Dynamics of Rank-1 Subspace
To characterize the evolution of the Rank-1 Subspace during training, we collect the sequence of $\bm{u}_{1}$ vectors across $T$ checkpoints for each module: $\mathcal{U}_{1}=\{\bm{u}_{1}^{(t)}\}_{t=1}^{T},$ which we refer to as the module’s update trajectory. Since each $\bm{u}_{1}^{(t)}$ lies in a high-dimensional space, we apply PCA for dimensionality reduction and then t-SNE for visualization. As shown in Figure 6 (a), the trajectories exhibit smooth, nearly linear patterns, with color gradients aligned to training progress, indicating the existence of a stable update direction. Limited by space, we describe the perspective and interpretation underlying this procedure in Appendix D.
Furthermore, to quantify whether there is a similar linear relationship between this evolution and reasoning performance, we treat each module’s Rank-1 trajectory $\bm{u}_{1}^{(t)}$ as the independent variable and the corresponding checkpoint’s accuracy $y$ on the reasoning dataset as the dependent variable, performing linear fitting using Partial Least Squares (PLS) regression (Geladi & Kowalski, 1986) and using $R^{2}$ as a measure of linearity (calculation details can be found in Appendix D). As shown in Figure 6 (b), some modules even achieve $R^{2}$ values close to 1, indicating that the Rank-1 update directions are strongly correlated with reasoning performance and can be effectively modeled by a fixed linear relationship. These experiments reveal the unique role of Rank-1 Subspace, where it acts as a bridge during training, providing the foundation for the visual correlation between training steps and reasoning performance.
@@ 3.2 Rank-1 Linearity & Module Importance
To address the first question, we aggregate $R^{2}$ values across all modules and layers in Appendix  14. Results show that MLP modules, particularly those in mid-to-high layers, tend to achieve higher $R^{2}$. A possible reason is that higher-layer modules are closer to the source of the reward signal, allowing them to better retain and utilize reasoning-related update directions. In contrast, self-attention modules generally exhibit lower $R^{2}$, suggesting noisier or partially redundant update signals.
@@ 4.2 Main Results
Table 1 presents the reasoning performance across six reasoning benchmarks (AIME24, AIME25, MATH, MINERVA, GPQA, and GSM8K) at different training stages, with and without AlphaRL-enhanced updates. The results show that AlphaRL significantly improves the model’s reasoning performance even at early stages (only 10% of the total training steps), achieving performance comparable to the non-AlphaRL model at 40% of training steps. By the 40% training stage, the AlphaRL-enhanced models almost reach the performance of fully trained models. For instance, RLOO and GRPO models achieve 96% of the reasoning performance of the fully trained model, surpassing the fully trained model on the MATH-500 dataset. For DAPO, AlphaRL reaches a relative accuracy of 102% on the GPQA dataset. Furthermore, we present an example after acceleration in Table 2, which shows that the AlphaRL-enhanced model not only remains consistent with the Fully Trained Model in evaluation metrics but also exhibits largely similar patterns and reasoning approaches in the responses.
Overall, the AlphaRL-enhanced models exhibit significant improvements at all training stages. AlphaRL not only accelerates the training process but also maintains reasoning performance close to that of the fully trained model while reducing training time. On average, AlphaRL accelerates the training process by up to 2.5 times while retaining the vast majority of the reasoning performance, demonstrating its tremendous potential in improving both training efficiency and reasoning capability.
@@ 5 Related Work
Reinforcement Learning for LLMs. Before the emergence of reasoning-capable models such as OpenAI’s o1, reinforcement learning (RL) was primarily employed in RLHF to improve instruction-following and alignment with human preferences (Ouyang et al., 2022). More recently, RL with Verifiable Rewards (RLVR) has been proposed as an effective strategy to enhance reasoning in domains such as mathematics and programming (Lambert et al., 2025). OpenAI’s o1 was the first to demonstrate that RL can incentivize large-scale reasoning, inspiring subsequent models such as DeepSeek-R1 (DeepSeek-AI et al., 2025), Kimi-K2 (Team et al., 2025), and Qwen3 (Yang et al., 2025). Among these, DeepSeek-R1 stands out for achieving strong reasoning capabilities via the online RL algorithm GRPO and for introducing the “Zero RL” paradigm, showing that reasoning can emerge even without conventional RL fine-tuning. Building on these advances, later approaches such as DAPO (Yu et al., 2025), VAPO (Yue et al., 2025b), and GSPO (Zheng et al., 2025) have further broadened the landscape of RL-based reasoning.
Emergent Behaviors of Reinforcement Learning. Yue et al. (2025a) investigated the differences in sampling between base models and RL-fine-tuned models, showing that RL improves sampling efficiency for pass@1 but does not directly enhance reasoning ability. Cui et al. (2025) identified the phenomenon of “entropy collapse” in reinforcement learning, where rapid early convergence causes the model to become overly confident, prematurely degrading its exploratory capacity. Wang et al. (2025a) observed in chain-of-thought reasoning that high-entropy tokens often act as branching points defining multiple potential reasoning paths. Shenfeld et al. (2025) compared RL with supervised fine-tuning and found that RL better preserves the model’s original knowledge and capabilities. Finally, Feng et al. (2025) demonstrated in an intent detection task that RL significantly improves generalization compared to SFT, underscoring the value of RL in more challenging scenarios.
@@ 7 Conclusion
Appendix A Experimental Setup

### v3

@@ Abstract
Recent advances in reasoning capabilities of large language models (LLMs) are largely driven by reinforcement learning (RL), yet the underlying parameter dynamics during RL training remain poorly understood. This work identifies two fundamental properties of RL-induced parameter updates in LLMs: (1) Rank-1 Dominance, where the top singular subspace of the parameter update matrix nearly fully determines reasoning improvements, recovering over 99% of performance gains; and (2) Rank-1 Linear Dynamics, where this dominant subspace evolves linearly throughout training, enabling accurate prediction from early checkpoints. Extensive experiments across 8 LLMs and 7 algorithms validate the generalizability of these properties. More importantly, based on these findings, we propose AlphaRL, a plug-in acceleration framework that extrapolates the final parameter update using a short early training window, achieving up to 2.5× speedup while retaining >96% of reasoning performance without extra modules or hyperparameter tuning. This positions our finding as a versatile and practical tool for large-scale RL, opening a path toward principled, interpretable, and efficient training paradigm for LLMs. Our model and code will be available at: https://github.com/caiyuchenustc/Alpha-RL.
@@ 1 Introduction
While effective, these researches mainly focus on post-hoc interpretability (i.e., explaining the endpoints of training), leaving the RL process itself less explored (Wang et al., 2025b; Zhang et al., 2025; Feng et al., 2025). Understanding parameter dynamics during RL is essential: not only for optimizing RL paradigms, but also for shedding light on the emergence of reasoning capabilities. Hence, this work aims to reveal the black-box of RL process by addressing two fundamental questions: Are RL-guided parameter updates governed by consistent principles, and how do these principles give rise to reasoning capabilities?
@@
To solve these, we conduct a step-wise analysis of the parameter update matrix $\Delta\bm{W}$ (i.e., the parameter difference between the RL-trained model and the base model). After applying mathematical tools such as orthogonal subspace projection (Cai and Cao, 2024), we uncover a striking phenomenon: performing singular value decomposition (SVD) (Koren et al., 2009) on $\Delta\bm{W}$ reveals that the top singular subspace, defined as the Rank-1 Subspace, almost entirely determines the reasoning gains from RL. That is, adding only the Rank-1 component of $\Delta\bm{W}$ to the base model is sufficient to recover nearly all of the RL-trained model’s reasoning improvements. More remarkably, this property holds not just at convergence, but at any intermediate step of RL training, as shown in Figure 2 (a). We formalize this finding as Property 1 (Rank-1 Dominance): Rank-1 Subspace of $\Delta\bm{W}$ determines the reasoning gains of LLMs throughout the RL training process.
@@
This property inspired us to probe how the Rank-1 Subspace evolves during RL training. By applying partial least squares (PLS) to track the dimension-wise trajectory of the Rank-1 Subspace across training steps, we observe an almost strictly linear upward trend, with the linear rate metric $R^{2}$ (Geladi and Kowalski, 1986) exceeding 0.96. Consequently, the Rank-1 Subspace at a target step can be accurately forecast from a short early window of training. We formalize this as Property 2 (Rank-1 Linear Dynamics): Rank-1 Subspace evolves approximately linearly with RL training process, yielding high predictability from early-stage checkpoints of RL process.
@@
To validate the generalizability, we conducted extensive experiments across 13 diverse LLMs (ranging from 7B to 32B parameters, e.g., Qwen3 (Yang et al., 2025), Llama3 (Grattafiori et al., 2024), and GLM4 (GLM et al., 2024)) trained with 10 advanced training algorithms (e.g., GRPO (Yu et al., 2025), Dr.GRPO (Liu et al., 2025b), and DAPO (Yu et al., 2025)). Our analysis shows that, for Property 1, the Rank-1 Subspace alone recovers an average of 99.17% of the reasoning capability. For Property 2, (1) the linearity of Rank-1 Subspace’s evolution exhibits an average $R^{2}$ of 0.914, and (2) predictions of its later state based on early-stage states achieve an average error of less than 5%. Crucially, control experiments with alternative training paradigms like supervised fine-tuning and distillation on the same models yielded neither property, demonstrating that these phenomena are distinctive characteristics of the RL process for LLMs. Detailed experimental setups and results are presented in Sections 2 and 3.
@@ 2 Dominance of Rank-1 Subspace (Property 1)
In this section, our objective is to analyze the effect of the Rank-1 Subspace of the parameter update matrix $\Delta W$ on reasoning gains. In Section 2.1, we first introduce the method for quantifying the contribution of the Rank-1 Subspace. Then, we exhibit experimental setups and main results in Section 2.2. Furthermore, the underlying causes of Rank-1 dominance and deeper analysis are investigated in Section 2.3.
@@ 2.1 Rank-1 Subspace
We first describe the method for quantifying the contribution of the Rank-1 Subspace to reasoning gains of RL training process. Specifically, performing SVD on $\Delta\bm{W}$, we have:
@@ 2.2 Rank-1 Subspace & Reasoning gains
Experiment Seting. Our experiments are conducted on eight models, covering five RL algorithms, including PPO (Schulman et al., 2017), RLOO (Ahmadian et al., 2024), GRPO (DeepSeek-AI et al., 2025), Dr.GRPO (Liu et al., 2025b), and DAPO (Yu et al., 2025), as well as Distillation (DIST) (Hinton et al., 2015) and Supervised Fine-tuning (SFT) (Ye et al., 2025). We comprehensively evaluate the reasoning performance of these models on six reasoning benchmarks, including AIME24, AIME25 (Ye et al., 2025) and MATH-500 (Lightman et al., 2023), to verify the robustness and generality of our findings. More detailed settings are provided in Appendix B.
@@ 2.3 Analysis of Rank-1 Dominance
Approximate Low-rank of $\Delta\bm{W}$. It is worth mentioning that, in the above experiments, we discovered a universal approximate low-rank (Zhang, 2015) property of $\Delta\bm{W}$ in RL, which is completely absent in SFT and DIST. Due to space limitations, we provide a detailed discussion of this phenomenon in Appendix C. Furthermore, we propose that the superior properties observed in RL-trained LLMs (e.g., minimizing catastrophic forgetting and improving generalization) may fundamentally arise from this low-rank structure, which plays a pivotal role in the model’s ability to effectively retain and adapt learned knowledge. Additionally, we also observed in our experiments the unique role of Rank-1 in guiding the reasoning chain, where modifying a small number of tokens achieves reasoning performance comparable to that of the fully trained model. We recommend interested readers to refer to the detailed results and analysis in Appendix D.
@@ 3.1 Exploring the Dynamics of Rank-1 Subspace
To characterize the evolution of the Rank-1 Subspace during training, we collect the sequence of $\bm{u}_{1}$ vectors across $T$ checkpoints for each module: $\mathcal{U}_{1}=\{\bm{u}_{1}^{(t)}\}_{t=1}^{T},$ which we refer to as the module’s update trajectory. Since each $\bm{u}_{1}^{(t)}$ lies in a high-dimensional space, we apply PCA for dimensionality reduction and then t-SNE for visualization. As shown in Figure 6 (a), the trajectories exhibit smooth, nearly linear patterns, with color gradients aligned to training progress, indicating the existence of a stable update direction. Limited by space, we describe the perspective and interpretation underlying this procedure in Appendix E.
Furthermore, to quantify whether there is a similar linear relationship between this evolution and reasoning performance, we treat each module’s Rank-1 trajectory $\bm{u}_{1}^{(t)}$ as the independent variable and the corresponding checkpoint’s accuracy $y$ on the reasoning dataset as the dependent variable, performing linear fitting using Partial Least Squares (PLS) regression (Geladi and Kowalski, 1986) and using $R^{2}$ as a measure of linearity (calculation details can be found in Appendix E). As shown in Figure 6 (b), some modules even achieve $R^{2}$ values close to 1, indicating that the Rank-1 update directions are strongly correlated with reasoning performance and can be effectively modeled by a fixed linear relationship. These experiments reveal the unique role of Rank-1 Subspace, where it acts as a bridge during training, providing the foundation for the visual correlation between training steps and reasoning performance.
@@ 3.2 Rank-1 Linearity & Module Importance
To address the first question, we aggregate $R^{2}$ values across all modules and layers in Appendix  15. Results show that MLP modules, particularly those in mid-to-high layers, tend to achieve higher $R^{2}$. A possible reason is that higher-layer modules are closer to the source of the reward signal, allowing them to better retain and utilize reasoning-related update directions. In contrast, self-attention modules generally exhibit lower $R^{2}$, suggesting noisier or partially redundant update signals.
@@ 4.2 Main Results
Table 1 presents the reasoning performance across six reasoning benchmarks (i.e., AIME24, AIME25, MATH, MINERVA, GPQA, and GSM8K) at different training stages, with and without AlphaRL-enhanced updates. The results show that AlphaRL significantly improves the model’s reasoning performance even at early stages (only 10% of the total training steps), achieving performance comparable to the non-AlphaRL model at 40% of training steps. By the 40% training stage, the AlphaRL-enhanced models almost reach the performance of fully trained models. For instance, RLOO and GRPO models achieve 96% of the reasoning performance of the fully trained model, surpassing the fully trained model on the MATH-500 dataset. For DAPO, AlphaRL reaches a relative accuracy of 102% on the GPQA dataset. Furthermore, we present an example after acceleration in Table 2, which shows that the AlphaRL-enhanced model not only remains consistent with the Fully Trained Model in evaluation metrics but also exhibits largely similar patterns and reasoning approaches in the responses. We present additional experimental results for more models in Appendix A.
Overall, the AlphaRL-enhanced LLMs exhibit significant improvements at all training stages. AlphaRL not only accelerates the training process but also maintains reasoning performance close to that of the fully trained model. On average, AlphaRL accelerates the training process by up to 2.5 times while retaining the vast majority of the reasoning performance, demonstrating its tremendous potential in improving both training efficiency and reasoning capability.
@@ 5 Related Work
Reinforcement Learning for LLMs. Before the emergence of reasoning-capable models such as OpenAI’s o1, RL was primarily employed in RLHF to improve instruction-following and alignment with human preferences (Ouyang et al., 2022). More recently, RL with Verifiable Rewards (RLVR) has been proposed as an effective strategy to enhance reasoning in domains such as mathematics and programming (Lambert et al., 2025). OpenAI’s o1 was the first to demonstrate that RL can incentivize large-scale reasoning, inspiring subsequent models such as DeepSeek-R1 (DeepSeek-AI et al., 2025), Kimi-K2 (Team et al., 2025), and Qwen3 (Yang et al., 2025). Among these, DeepSeek-R1 stands out for achieving strong reasoning capabilities via the online RL algorithm GRPO and for introducing the “Zero RL” paradigm, showing that reasoning can emerge even without conventional RL fine-tuning. Building on these advances, later approaches, such as DAPO (Yu et al., 2025), VAPO (Yue et al., 2025b), GSPO (Zheng et al., 2025) and CISPO (MiniMax et al., 2025), have further broadened the landscape of RL-based reasoning.
Emergent Behaviors of Reinforcement Learning. Yue et al. (2025a) investigated the differences in sampling between base models and RL-fine-tuned models, showing that RL improves sampling efficiency for pass@1 but does not directly enhance reasoning ability. Cui et al. (2025) identified the phenomenon of “entropy collapse” in reinforcement learning, where rapid early convergence causes the model to become overly confident, prematurely degrading its exploratory capacity. Wang et al. (2025a) observed in chain-of-thought reasoning that high-entropy tokens often act as branching points defining multiple potential reasoning paths. Shenfeld et al. (2025) compared RL with supervised fine-tuning and found that RL better preserves the model’s original knowledge and capabilities. Finally, Feng et al. (2025) demonstrated in an intent detection task that RL significantly improves generalization compared to SFT, underscoring the value of RL in more challenging scenarios, while Mukherjee et al. (2025) identified the sparse nature inherent in RL.
@@ 7 Conclusion
8 Acknowledge
The Major Project of the National Social Science Fund of China: “Key Enabling Technologies for Agentic Document and Information Services in the Context of Digital and Intelligent Transformation” (Project No. 23&ZD228)
Appendix A Additional Experiment
Table 3: AlphaRL’s performance on Qwen3-8B-Base, Qwen3-14B-Base and GLM-4-9B-0414.
Stage AIME24 AIME25 MATH MINERVA GPQA GSM8K Avg.
Qwen3-8B-Base
Training 5% 7.50 5.50 62.25 27.25 31.50 78.47 35.45
Training 5%+AlphaRL 10.00 6.67 69.75 31.15 34.75 81.65 38.91
Training 10% 12.50 7.50 70.25 32.07 36.66 84.30 40.55
Training 10%+AlphaRL 15.00 11.67 76.45 40.46 41.54 93.75 46.47
Training 20% 13.17 7.50 72.25 35.07 38.50 87.54 42.34
Training 20%+AlphaRL 16.67 12.50 75.00 41.25 42.85 94.35 47.10
Training 30% 14.45 11.67 74.25 36.37 39.42 91.25 44.56
Training 30%+AlphaRL 23.33 20.00 78.75 41.85 44.13 93.20 50.21
Training 40% 15.80 11.67 77.60 37.07 41.67 93.20 46.30
Training 40%+AlphaRL 28.33 23.75 80.50 43.27 49.25 94.75 53.31
Training 50% 23.33 20.00 78.50 39.87 44.33 93.75 49.40
Training 50%+AlphaRL 27.67 23.33 82.00 42.85 47.75 94.50 53.01
Fully Trained Model 28.54 24.17 80.95 44.02 48.23 94.35 53.38
Qwen3-14B-Base
Training 5% 12.50 8.33 74.30 40.44 41.04 90.39 44.50
Training 5%+AlphaRL 16.11 11.11 77.75 42.27 43.25 91.25 46.96
Training 10% 13.67 10.17 77.50 41.58 43.50 91.50 46.99
Training 10%+AlphaRL 21.72 15.05 81.25 42.37 44.76 92.75 49.98
Training 20% 22.00 15.33 81.25 42.37 45.67 92.75 49.89
Training 20%+AlphaRL 25.28 21.11 86.25 44.65 48.76 94.75 53.47
Training 30% 23.33 18.33 85.00 43.67 47.34 93.50 51.53
Training 30%+AlphaRL 30.69 23.89 89.50 46.50 51.73 96.75 56.84
Training 40% 28.33 20.00 88.25 44.30 49.75 94.75 54.56
Training 40%+AlphaRL 38.47 31.25 91.25 47.75 52.59 97.75 59.18
Training 50% 37.64 28.89 89.75 45.67 51.25 95.25 58.08
Training 50%+AlphaRL 40.00 31.80 92.25 48.33 53.75 97.50 60.27
Fully Trained Model 40.50 32.63 91.75 48.33 54.50 97.50 60.87
GLM-4-9B-0414
Training 5% 4.17 1.67 64.20 35.29 40.40 87.34 38.84
Training 5%+AlphaRL 5.67 3.33 66.75 37.35 42.50 88.85 40.91
Training 10% 12.50 7.50 72.25 37.25 43.50 88.50 43.58
Training 10%+AlphaRL 16.67 11.67 79.50 40.25 53.25 90.75 48.68
Training 20% 15.00 13.33 79.50 39.75 44.50 90.25 47.06
Training 20%+AlphaRL 20.33 17.00 83.25 43.50 57.75 92.25 52.01
Training 30% 18.52 15.06 84.20 42.33 46.36 92.33 50.13
Training 30%+AlphaRL 24.54 19.33 88.50 48.25 53.33 95.55 54.92
Training 40% 22.67 21.50 86.50 45.50 49.35 94.00 53.92
Training 40%+AlphaRL 29.33 27.25 91.50 50.25 54.75 96.25 58.89
Training 50% 24.33 22.67 87.25 46.75 50.20 94.25 54.57
Training 50%+AlphaRL 30.60 30.00 91.50 50.87 55.33 97.00 59.72
Fully Trained Model 31.10 29.75 91.75 51.34 55.67 96.25 59.98
Table 4: AlphaRL’s performance on Deepseek-R1-Distill-Qwen-7B and Llama3-8B-Instruct.
Stage AIME24 AIME25 MATH MINERVA GPQA GSM8K Avg.
Deepseek-R1-Distill-Qwen-7B (After Distillation)
Training 5% 50.83 39.17 91.25 50.00 38.89 93.00 60.69
Training 5%+AlphaRL 51.92 40.08 92.75 51.35 40.75 93.00 61.64
Training 10% 52.50 41.75 92.75 51.50 41.17 94.25 62.32
Training 10%+AlphaRL 56.50 44.50 94.50 54.25 43.85 94.75 64.73
Training 20% 56.50 45.75 93.75 53.75 44.25 94.75 64.79
Training 20%+AlphaRL 60.00 52.25 95.50 57.85 48.25 95.50 48.23
Training 30% 59.50 51.50 95.20 56.22 47.25 95.50 67.52
Training 30%+AlphaRL 64.50 57.25 96.17 60.07 51.33 96.50 70.97
Training 40% 63.25 55.25 95.00 58.45 49.86 96.50 69.74
Training 40%+AlphaRL 67.50 61.50 97.25 62.35 53.67 97.75 73.34
Training 50% 65.75 58.85 96.50 60.65 52.27 97.25 71.87
Training 50%+AlphaRL 68.25 62.50 97.25 62.58 54.16 97.75 73.75
Fully Trained Model 68.75 62.00 97.50 61.74 54.16 98.25 73.65
Llama3-8B-Instruct (After SFT)
Training 5% 9.58 6.67 68.75 26.44 29.75 85.25 37.74
Training 5%+AlphaRL 12.36 9.02 71.25 28.37 31.25 86.50 39.79
Training 10% 12.63 9.16 73.00 28.37 31.75 86.25 40.19
Training 10%+AlphaRL 15.97 13.06 77.75 31.76 34.54 89.25 43.72
Training 20% 20.00 15.00 78.00 31.07 34.75 89.50 45.06
Training 20%+AlphaRL 22.67 20.00 84.75 34.75 37.25 92.25 48.94
Training 30% 23.67 20.33 81.50 35.25 38.25 91.50 48.75
Training 30%+AlphaRL 30.64 27.02 85.75 38.85 42.13 93.50 53.98
Training 40% 28.33 23.67 84.25 37.50 40.75 92.25 51.46
Training 40%+AlphaRL 40.75 30.00 87.50 41.50 44.25 94.50 56.75
Training 50% 31.11 27.11 86.25 37.50 42.25 92.50 52.79
Training 50%+AlphaRL 39.12 30.75 87.75 42.25 43.75 94.25 56.98
Fully Trained Model 38.75 31.25 87.50 42.75 43.75 95.00 56.83
In Table 3, we report the full performance trajectories of the 9B and 14B models across training stages to demonstrate that the observed patterns persist at larger scales. As shown, AlphaRL exhibits similarly stable extrapolation behavior on these larger models, accurately reproducing the performance gains associated with later-stage updates even when applied at early checkpoints. Taken together, these results further support a central conclusion of this work: the predictability of RL dynamics is not limited to mid-sized models, but continues to hold as model scale increases.
Table LABEL:appendax4 extends our analysis to a broader set of base models, including distilled variants and models that have undergone instruction tuning on mathematics-focused datasets. Despite these differences in pretraining objectives and data distributions, the trends observed in the main experiments remain consistent. In particular, both the Rank-1 structure of RL updates and the effectiveness of AlphaRL hold across these diverse model initializations. This consistency suggests that the predictable dynamics identified in this work are not tied to a specific base model configuration, but instead reflect a general property of RL training that persists under architectural variations.
Table LABEL:appendax5 and LABEL:appendax6 further extends our analysis to an adversarial self-play setting, allowing us to examine whether the regularities identified in this work persist beyond traditional mathematical and reasoning tasks. Specifically, we adopt the Self-Play framework proposed by Liu et al. (2025a), which is conceptually similar to AlphaZero (Silver et al., 2017) in Go and enables a single language model to play both competing roles within the same training process. In this setup, we use Qwen-4B-Base as the base model, train it for roughly 200 checkpoints (saving one checkpoint every two epochs), and evaluate its win rate against Gemini-2.0-Flash-Lite3 throughout training.
Despite the substantial differences from mathematical reasoning—such as the adversarial nature of the environment, the discrete and game-specific dynamics, and the inherently noisy win/loss reward signal—the results exhibit the same trends observed in our main experiments: RL updates consistently maintain a stable Rank-1 structure, and AlphaRL accurately extrapolates late-stage performance improvements using only early-stage checkpoints. These findings indicate that the predictive structure uncovered in this work is not tied to any particular task formulation. Instead, it generalizes beyond reasoning tasks and remains valid even in challenging RL settings characterized by high-variance rewards and strong adversarial interactions.
Table 5: AlphaRL’s performance on Kuhn Poker Games.
Stage WinRate Rank-1 WinRate AlphaRL WinRate
Qwen3-4B-Base
Training 5% 3% 2% 4%
Training 10% 5% 5% 12%
Training 20% 13% 15% 28%
Training 30% 27% 25% 41%
Training 40% 43% 44% 59%
Training 60% 52% 52% 64%
Training 80% 61% 59% -
Training 100% 65% 62% -
Table 6: AlphaRL’s performance on MATH-500 (after training on Kuhn Poker Games).
Stage Full Model Rank-1 AlphaRL
Qwen3-4B-Base
Training 5% 45% 45% 51%
Training 10% 52% 51% 60%
Training 20% 56% 58% 65%
Training 30% 61% 61% 69%
Training 40% 64% 64% 75%
Training 60% 68% 68% 81%
Training 80% 73% 71% -
Training 100% 77% 77% -
Table 7: L2 norm of updates across methods and the fraction of update information captured by the unscaled Rank-1 and Rank-1% Subspaces.
Method Offline DIST Online DIST Offline SFT Offline DPO Online RL DAPO
On-Policy Sampling No Yes No No Yes
Constraints on the distribution of $\pi_{\theta}$ No Yes No Yes Yes
Average Update Norm 21.24 1.46 10.75 0.79 0.01
Rank-1 fraction 12.6% 10.45% 7.7% 11.42% 19.3%
Rank-1% fraction 29.3% 24.78% 27.7% 24.13% 36.5%
Table 8: Performance under Rank-1 and Rank-k% Subspaces on MATH-500.
Method Offline DIST Online DIST Offline SFT Offline DPO Online RL DAPO
Rank-1 1.50% 56.40% 2.50% 62.25% 100.00%
Rank-1% 2.50% 64.50% 3.00% 71.50% 100.00%
Rank-10% 3.00% 71.50% 4.00% 74.50% 100.00%
Rank-20% 3.00% 72.50% 7.50% 79.00% 100.00%
Rank-30% 3.50% 77.00% 8.00% 83.25% 100.00%
Rank-50% 3.50% 80.25% 12.25% 88.25% 100.00%
Rank-70% 16.00% 84.00% 24.00% 88.25% 100.00%
Rank-90% 23.25% 89.25% 46.75% 93.25% 100.00%
Appendix B Experimental Setup